### v3

@@ 6.3 Active learning sample complexity
According to the Blumer bound blumer_occams_1987 , for a version space learner that randomly samples examples to produce a hypothesis of error at most $\epsilon$ with a probability less than or equal to $\delta$, the sample complexity is $s_{passive}>\frac{1}{\epsilon}(ln(|H|)+ln(\frac{1}{\delta}))$. Following this, it must be true that $\epsilon\,s_{passive}>(\epsilon+\phi)\,s_{active}$. So we can derive the Inequality 2 in the theorem:
@@
This theorem shows the sample complexity gain of active learning over learning from randomly sampled instances. Theorem 6.1 says that the number of instances needed to select by active learning should be some factor smaller than the number of randomly sampled examples given some predictive accuracy level. Given a desirable predictive error $\epsilon$, this factor can be computed if the minimal reduction ratio $\phi$ can be estimated from the instance space. When selected instances have a larger minimal reduction ratio, the sample complexity gain would be more significant. On the other hand, if instances have low discriminative power and the minimal reduction ratio tends to zero, the sample complexity gain would be minimal. In Section 8, we empirically show empirical evidence supporting this theorem for $BMLP_{active}$ and randomly sampled experiments when learning gene functions in a GEM.
7 Implementation
7.1 Encoding a metabolic network
Our encoding of a metabolic network is based on the Robot Scientist King04:RobotScientist ; ase_progol . The Robot Scientist is implemented using the ILP system Progol progol and finds the synthesisable set of metabolites by traversing a metabolic network. It uses a metabolic network as a deductive database and answers a query of user-defined gene deletion and experimental conditions based on SLD resolutions. Its deductive database is written in Prolog syntax and uses Prolog list operations to manipulate constant symbols representing metabolic substances. Metabolic reactions are enumerated based on the set of substances currently synthesisable. The products of a metabolic reaction extend the list of synthesisable substances. Since the number of metabolic reactions is finite, the closure of members in this list needs to contain essential substances for the growth of an organism. In an auxotrophic mutant experiment outcome prediction, a gene function is removed. Some key metabolic reaction pathways might no longer be viable due to the removal, leading to insufficient substances for growth. In this case, the predicted experiment’s outcome is a lack of growth and a phenotypic effect.
@@ 7.1 Encoding a metabolic network
In the following program $\mathcal{P}_{2}$, the $codes/2$ clause says that the gene locus “b2379” is responsible for the enzyme “enz201”. The $enzyme/6$ clause defines the reversible reaction “alata_l” (L-alanine transaminase). This reaction involves reactants “akg_c” (Oxoglutarate), “ala__L_c” (L-alanine) and products “glu__L_c” (L-glutamate), “pyr_c” (Pyruvate). An integer value in {1,2} indicates the reaction reversibility and the integer “2” says that the reaction is reversible.
$\mathcal{P}_{2}:\left\{\begin{aligned} codes(b2379,enz201).&\\ enzyme(enz201,[b2379],[alata\_l],2,[[akg\_c,ala\_\_L\_c]],&\\ [[glu\_\_L\_c,pyr\_c]]).&\end{aligned}\right\}$
@@
Our encoding of the GEM iML1515 also contains $codes/2$ and $enzyme/6$ definitions. These clauses enable us to describe metabolic reactions via boolean matrices. We convert gene-reaction-metabolite mappings and standard identifiers in iML1515111The iML1515 model is described in the text-based JavaScript Object Notation (JSON). It uses standard identifiers from the Biochemical Genetic and Genomic (BiGG) king_bigg_2016 knowledge base of large-scale metabolic reconstructions. into a logic program containing $codes/2$ and $enzyme/6$ clauses. iML1515 uses standard nomenclature so the logic program encoding can be applied to other genome-scale metabolic networks. The logic program encoding can also be linked to publications and external databases to obtain further details on genes, proteins, reactions and metabolites.
7.2 Simulating a GEM using BMLP-IE
@@
Another difference is that boolean matrices are created using the $reaction/2$ clauses for bottom-up evaluation. Algorithm 2 shows the application of BMLP-IE (Algorithm 1) for finding the transitive closure of synthesisable substances. The two boolean matrices inputs to the BMLP-IE algorithm represent reactants and products participating in the reactions. The growth medium condition is encoded as the input vector to the BMLP-IE algorithm. This vector would be extended to represent substances that are synthesisable by the organism in an auxotrophic mutant experiment. Specifying knockout genes excludes a subset of reactions. A vector computed during an intermediate step in the BMLP-IE algorithm denotes viable reactions given the set of knockout genes. A viable reaction is indicated by a binary bit 1 in this vector. Blocked reactions are excluded based on $codes/2$ and $enzyme/6$ definitions. In BMLP-IE, boolean operations compute the transition between metabolic states and return the closure set of synthesisable substances. A binary phenotypic classification is made according to the inclusion of essential substances in the closure set, indicating cell growth comparable to an unedited wild-type MG1655 E. coli strain. A binary classification of “1” corresponds to a phenotypic effect and “0” is no phenotypic effect. The essential substances come from the biomass function of the wild-type in iML1515 iML1515  [^2] We use the wild-type biomass function specified by iML1515 from http://bigg.ucsd.edu/ king_bigg_2016 . We use the qualitative version of this biomass function, meaning that we only care about the substances in the biomass, not their reaction coefficients.. The biomass function describes the necessary substance composition for the growth of the organism.
7.3 Active learning using BMLP
Algorithm 3 $BMLP_{active}$
Output: A hypothesis $h_{max}\in H$.
1:  Let $T_{selected}=\emptyset$, $V_{0}=H$.
2:  while $cost(T_{selected})\leq C$ and $|V_{j}|>1$ and $|T_{selected}|<N$ do
3:     if $T_{selected}==\emptyset$ then
4:        Select $t_{j}$ with the largest minimum reduction ratio $p(t_{j},V_{j})$
5:           from remaining experiments in $T$ to add to $T_{selected}$.
6:     else
8:     end if
9:     break if $cost(T_{selected})>C$
10:     Consult the label $l$ of $t_{j}$ from the oracle $O$.
12:     $V_{j+1}=\{h_{i}\,|\,h_{i}\in V_{j}\text{ and }\textbf{R}_{i,j}=l\}$.
13:     For $h_{i}\in V_{s+1}$, compute $compression(h_{i},T_{selected})$.
14:     $h_{max}$ is the hypothesis with the largest compression score.
15:  end while
Active learning in $BMLP_{active}$ is described by Algorithm 3. The input background knowledge is created from a GEM model. The user needs to provide a set of candidate experiments, candidate hypotheses and an experiment cost function. Labels of experiment instances are requested from a data source, e.g. a laboratory or an online dataset. Labelled experimental data are considered ground truths. Initially, we randomly shuffle all candidate experiments and select an experiment with the most discriminative power if its cost is under budget. Alternatively, a discriminative experiment with the lowest cost can be selected for a fixed initialisation.
In each active learning cycle in $BMLP_{active}$, hypotheses inconsistent with ground truth experimental outcomes are pruned. $BMLP_{active}$ selects experiments to minimise the expected value of a user-defined cost function (see Section 6). We apply BMLP-IE to obtain predicted experiment outcomes and compare them with experimental data. For a hypothesis space $H$ and a set of experiment instances $T$, BMLP-IE produces predicted outcomes in a boolean matrix R of size $|H|\times|T|$. Given the j-th actively selected instance $t_{j}\in T$ and $h_{i}\in V_{j}\subseteq H$, $(\textbf{R})_{i,j}=1$ if $h_{i}\cup BK\models t_{j}$ and otherwise $(\textbf{R})_{i,j}=0$. We refer to R as a classification table. In experiments from Section 8, we construct classification tables by predicting phenotypic effects for all combinations of candidate experiments and hypotheses. This avoids repeating simulations in hypothesis pruning and compression calculation. When computing hypothesis compression, we estimate the generality of a hypothesis by calculating its coverage for all unlabelled instances based on the classification table. Currently, we only consider labels from synthetic or online phenotype data. However, $BMLP_{active}$ can be coupled with a high-throughput experimental workflow to automate experiments.
8 Experiments
@@
Since a logic program representing a GEM describes many biological entities and intricate relations between them, abduction of gene function hypotheses and in-vivo validations requires exploring a large hypothesis space and instance space. Both are challenging computationally and empirically. $BMLP_{active}$ aims to address these by evaluating a large logic program with boolean matrices and actively selecting experiment instances to perform. In our experiments [^3] All experiments are performed on a workstation with Intel(R) Core(TM) i9-7900X CPU 3.30GHz (20 CPUs). , we intend to demonstrate that the $BMLP_{active}$ framework can efficiently perform phenotypic predictions and actively suggest informative experiment instances that can reduce experiment costs.
@@
In this work, we define the experimental cost function as the cost of the reagents which are optional nutrient substances used in selected experiments. We also assume when added to growth media, reagents are used in the same volume. In addition, we do not explicitly account for plastic consumables, equipment deprecation or specify the time to carry out each experiment, which we consider fixed and are affected mainly by the number of experiments. We also do not account for costs incurred due to different experiment types such as different measurement timescales that prevent experiments from being performed concurrently. To maintain consistent scaling for comparisons, the total experimental reagent cost is the amount of unit cost required by selected experiments. We consider the unit cost of the reagents as the cost of the cheapest reagent. We calculate the total experimental reagent cost as the sum of the reagent costs [^4] We provide the costs of reagents in Table 2 and 3 of Appendix A. in selected experiments divided by the unit cost.
Q1: Can BMLP-IE reduce the runtime of simulating GEM?
Using only Prolog to simulate metabolic networks requires grounding viable reaction pathways based on SLD resolutions. We aim to demonstrate that a general-purpose Prolog interpreter can be augmented with our BMLP approach for faster phenotypic simulations given logically encoded metabolic network models. $BMLP_{active}$ extends the logical representation of a metabolic network in the Robot Scientist King04:RobotScientist ; ase_progol by additionally applying boolean matrix computation via BMLP. Therefore, we consider the Robot Scientist’s logical encoding as a benchmark. Since the Robot Scientist was implemented using Progol progol , we adapt its metabolic network encoding to SWI-Prolog. To answer Q1, we compare the simulation runtime of $BMLP_{active}$, which uses SWI-Prolog and BMLP-IE, against the Robot Scientist’s encoding in SWI-Prolog.
Q2: Does BMLP-IE accurately predict phenotypes?
BMLP-IE makes binary phenotypic predictions and accounts for the necessary growth conditions of the organism. In contrast to qualitative predictions produced by BMLP-IE, FBA palsson_systems_2015 is more commonly used in the literature to quantitatively predict the cell growth rate based on fluxes of biomass metabolites at the steady state. We compare BMLP-IE phenotypic predictions against those given by FBA from the iML1515 model in iML1515 . To answer Q2, the comparison between BMLP-IE and FBA predictions is made with respect to ground truth experimental data also in iML1515 .
Q3: Can $BMLP_{active}$ reduce the cost for learning gene functions in GEM?
We consider two measures of experiment cost: the experimental resource cost and the number of experiments. These two costs are related since the number of experiments generally is the main driver of experimental cost. We aim to show that experiments selected by $BMLP_{active}$ reduce overall experimental reagent cost compared to random experiment selection. We additionally demonstrate that $BMLP_{active}$ lowers the number of experiments (the sample complexity) needed to learn accurate gene functions. Gene function recovery experiments are conducted to answer Q3. To demonstrate that $BMLP_{active}$ can recover known gene functions, we delete parts of iML1515 to create an incomplete knowledge base artificially. Active learning from the incomplete model is a mock of an actual discovery. Active learning aims to recover deleted gene functions within a reagent cost budget and with a limited number of experiments. We also examine Q3 for recovering isoenzyme functions which require navigating a large experimental design space.
Q4: Does Theorem 6.1 correctly describe the relationship between the sample complexity of $BMLP_{active}$ and random experiment selection?
Theorem 6.1 shows the advantage of active learning in improving the efficiency of experimentation. We empirically examine the sample complexity bound specified in Theorem 6.1. Let the sample complexity of $BMLP_{active}$ and a random experiment sampling learner be $s_{active}$ and $s_{passive}$ respectively. From Theorem 6.1 we can obtain the following inequality:
$\frac{\epsilon}{s_{active}}\geq\frac{\epsilon+\phi}{s_{passive}}+c\text{ with c as constant}$
In order to answer Q4, we use $BMLP_{active}$ and random experiment sampling to recover gene functions as described above. We calculate the ratios of the error rates of learned gene functions over the number of experiments selected.
Q5: Can $BMLP_{active}$ learn digenic interactions?
While isoenzyme functions are identified as one source of errors in iML1515 bernstein_evaluating_2023 , full exploration of the gene-isoenzyme relations is laborious. This motivates us to answer Q5. We perform gene function recovery experiments using $BMLP_{active}$ to rediscover a key isoenzyme for demonstration. We focus on the aromatic amino acid biosynthesis pathway. It is a critical pathway for micro-organism survival that contains important gene targets for engineering high aromatic amino acid production strains. Although this pathway has been intensively studied for decades, gaps remain in the understanding of gene-gene relations price_filling_2018 . We show that an isoenzyme function in the aromatic amino acid pathway can be accurately recovered via active learning. This demonstrates when $BMLP_{active}$ is integrated with an experimental workflow that systematically perturbs genes under different growth conditions, we can potentially reveal novel insights for metabolic engineering and strain optimisation.
8.1 Experiment 1: BMLP-IE simulations and predictions
Single thread (seconds) 20 CPUs (seconds)
SWI-Prolog $37.842\pm 9.668$ $5.089\pm 0.119$
SWI-Prolog + BMLP-IE $\mathbf{0.220\pm 0.099}$ $\mathbf{0.061\pm 0.008}$
Table 1: Simulation runtimes in CPU time. BMLP-IE leads to a 170 times improvement in simulation time efficiency. Multi-threading BMLP-IE enhances runtime efficiency 600 times compared to base SWI-Prolog.
@@ 8.1 Experiment 1: BMLP-IE simulations and predictions
Materials. In each simulation, we used $BMLP_{active}$ to predict a phenotype in an auxotrophic single-knockout mutant experiment. Every experiment removes one gene from the biological system and uses a growth medium. The GEM iML1515 has been validated against 16 carbon sources as optional nutrients in iML1515 . We focused on the same experimental conditions for simulations with base SWI-Prolog and $BMLP_{active}$. Experimental phenotypic data on the same 16 carbon sources and FBA phenotypic predictions of iML1515 are already processed as binary classes in iML1515 . In iML1515 , measured and predicted mutant growth rates at a proportion of the wild-type’s growth rate greater than a fixed threshold are mapped to the no phenotypic effect class.
@@
Methods. We randomly sampled 100 simulations to perform and compute the average wall time for each computation method. We first compared the mean and standard deviation in simulation time of our new logical matrix encoding against the Prolog encoding. In addition, we explored multi-threading with both encodings. We used the multi-threading library offered by SWI-Prolog. Owing to the overhead of managing threads, multi-threading boolean operations with SWI-Prolog have diminishing returns. In BMLP-IE, simulation parameters are encoded as matrices, enabling us to run simulations in batches. For BMLP-IE, we used the concurrent threading library in SWI-Prolog at the batch level. The runtime experiments with the multi-threading are repeated 10 times. We show this comparison in Table 1.
Figure 4: Normalised confusion matrices of (a) BMLP-IE and (b) FBA phenotypic predictions based on the GEM model iML1515. Each number has been normalised by the size of the experiment dataset in iML1515 . From this dataset, the phenotypic prediction accuracies of BMLP-IE and FBA are $88.7\%$ and $93.4\%$ respectively.
To evaluate BMLP-IE’s accuracy, we compared its phenotypic predictions against experimental phenotypic data on the same 16 carbon sources. We identified the number of consistent and inconsistent predictions and created normalised confusion matrices for BMLP-IE and the FBA predictions from iML1515 . We show these normalised confusion matrices in Fig 4.
Results and discussion. Table 1 shows a 170 times improvement in simulation time by BMLP-IE compared to base SWI-Prolog. The simulation runtime is further improved with multi-threading. With 20 cores, BMLP-IE uses only $\frac{1}{600}$ of the simulation time compared to running simulations with base SWI-Prolog. However, the multi-threading runtime improvement is not as pronounced as the runtime improvement from using BMLP-IE. This result answers Q1 as the SWI-Prolog implementation of BMLP-IE significantly improves the phenotype simulation runtime of iML1515. While BMLP-IE is a specifical-purpose algorithm, this result shows that a machine learning system implemented in SWI-Prolog can be more time-efficient if it uses BMLP-IE when the background knowledge can be described as a linear and immediately recursive datalog program.
In Fig 4, the confusion matrices show a comparable predictive accuracy by BMLP-IE ($88.7\%$) to that of standard FBA ($93.4\%$). This answers the question Q2. Compared with FBA’s growth rate predictions, the misclassifications by BMLP-IE are mainly due to it not being able to predict reduced cell growth. FBA permit quantitative descriptions of metabolic fluxes but metabolites are only binarily represented in an OEN. However, the strength of the BMLP-IE approach is that it is easily adaptable for extending the underlying GEM. FBA methods are mainly used as mathematical modelling tools so they are not immediately applicable for learning.
@@ 8.2 Experiment 2: active learning sample complexity and experiment cost
Materials. Up to 3 of 16 common carbon sources were used as additional nutrients to the base growth medium. Each of these 16 carbon sources is an optional nutrient in the experimental data for validating iML1515 . This set contains preferred carbon sources martinez-gomez_new_2012 ; zampieri_regulatory_2019 for E. coli such as glucose and glycerol. We first simulated iML1515 with up to 3 optional carbon sources, generating $\sum_{0}^{i=3}\binom{16}{i}=697$ synthetic data. This allows us to go beyond the 16 experimental data for each gene knockout in iML1515 . We observed a single-knockout phenotypic effect from 213 genes. We treated these synthetic data as ground truths where 213 genes have only positive examples. We only focused on a subset of 2 genes because they have both positive and negative examples [^5] The distribution of training examples for this experiment is described in Table 4 of Appendix B.. Then, we removed gene functions associated with these 2 genes from the background knowledge. Thus, the resulting incomplete model has a smaller set of $codes/2$ clauses.
@@
Methods. We focused on one removed gene function at a time to recover it. Each hypothesis except the empty hypothesis associates a gene locus identifier, whose function we tried to recover, with an $enzyme/6$ clause. We concentrate on 1294 $enzyme/6$ clauses that represent the non-maintenance reactions in iML1515. Therefore, the hypothesis space had 1295 $codes/2$ hypotheses, including an empty hypothesis. The instance space contained 697 experiments with optional growth nutrients up to 3 carbon sources.
Figure 5: Total experimentation cost needed for the recovered models to reach different levels of predictive accuracy. Models recovered from random experiment selection and $BMLP_{active}$ can both converge to 100% predictive accuracy. However, models recovered by $BMLP_{active}$ reach 100% predictive accuracy with $\frac{1}{10}$ reagent cost expenditure of random experiment sampling. The predictive accuracy of an empty hypothesis is 17.4%.
$BMLP_{active}$ actively selected experiments from the instance space to reduce the hypothesis space. The random selection method randomly sampled instances from the instance space. First, we constrained the selection of experiments by an experimental reagent cost budget $C$. After experiment selection, both methods returned the hypothesis with the single highest compression, otherwise, a hypothesis was selected from those sharing the highest compression randomly. The final hypotheses obtained by either method would output a recovered model. We compared $BMLP_{active}$ with the random experiment selection strategy with increasing experimental reagent cost budget $C\in\{1,10,50,100,500,1000,5000,10000\}$. We tested the predictive accuracy of the recovered models against all 697 synthetic data. The cost of selected experiments and predictive accuracy of the recovered model were recorded for $BMLP_{active}$ and random selection. Each selection method was repeated 10 times for each reagent cost budget $C$. We show this comparison in Fig 5.
Figure 6: Reduction of the number of experiments needed to recover gene functions (sample complexity reduction). $BMLP_{active}$ learns 100% accurate gene functions with 3 experiments while random sampling requires 25 experiments to do so. The predictive accuracy of an empty hypothesis is 17.4 %. Predictive accuracy is higher for both methods initially since experiment selection is no longer constrained by reagent costs.
In addition, we constrained the number of experiments $N$ that can be selected by increasing it linearly according to $\{0,1,2,3,4,5,10,15,20,25,30\}$. We applied $BMLP_{active}$ and random experiment selection with compression to decide the final hypothesis for recovering models. Each selection method was repeated 10 times for every number of selected experiments $N$. Then, we computed the phenotypic predictive accuracy of the recovered models against all 697 synthetic data. We show this comparison in Fig 6. For experiment number limits excluding 0, we calculated the mean error of learned hypotheses and divided it by the number of experiments selected to obtain the ratio. In Fig 7, we demonstrate how these ratios change to support Theorem 6.1.
Results and discussion. Fig 5 shows that $BMLP_{active}$ spends 10% of the cost used by random experiment sampling when converging to the correct hypothesis. The y-axis is the total experimental reagent cost in $log_{10}$ calculated from selected experiments. And 100% predictive accuracy indicates successful recovery of the deleted gene functions. The correct gene functions were recovered by $BMLP_{active}$ and random experiment selection with $10^{1.99}$ and $10^{2.99}$ experimental reagent costs respectively. The total experimental reagent cost is computed by dividing the unit cost which is the cost of the cheapest optional nutrient (£$\pounds 0.038/gram$ ). Given the current hypothesis space and experimental design space and assuming each reagent was used one gram, experiments selected by $BMLP_{active}$ and random sampling would cost $10^{1.99}\times 0.038\approx\pounds 3.8$ and $10^{2.99}\times 0.038\approx\pounds 38$ to learn a single gene function. Consideration of differential resource factors in the user-defined experimental cost function could significantly increase the total experimental cost.
We then examined the advantage of $BMLP_{active}$ over random experiment selection on the number of experiments needed to recover gene functions. The number of experiments is a significant determinant of experimental cost. Fig 6 shows that $BMLP_{active}$ reduces the number of instances needed for labelling to learn accurate gene functions. While random experiment selection requires 25 experiments to recover the deleted gene functions, $BMLP_{active}$ only needed 3 experiments. Fig 5 and Fig 6 answer the question Q3 since $BMLP_{active}$ successfully reduced the experimental reagent cost and experiment cost compared to random experiment selection. Fig 5 and Fig 6 show $BMLP_{active}$ can simultaneously optimise a user-defined experimental cost function and reduce the total number of experiments to learn gene functions accurately. This means that $BMLP_{active}$ is highly flexible and could be tailored for specific experimental objectives in a discovery process. Later, in Experiment 3, we revisit the question Q3 for recovering a known gene-isoenzyme association which presents a bigger empirical challenge due to the need to explore a larger experimental design space.
Fig 7 demonstrates the mathematical relationships between predictive error and the number of experiments required by $BMLP_{active}$ and random experiment selection to converge to the correct hypothesis. For this gene function recovery task, $BMLP_{active}$ selected experiments with minimal reduction ratio $\phi\geq 0.5\%$. The ratio $\frac{\epsilon}{s_{active}}$ is greater than $\frac{\epsilon+\phi}{s_{passive}}$ for error $\epsilon>0$. There are fewer points on the active learning ratio curve since $BMLP_{active}$ converges to 100% accuracy using fewer training examples. The two curves intersect at $\epsilon=0$ since both converged to 100% accuracy. Fig 7 answers the question Q4 and supports Theorem 6.1. Theorem 6.1 shows that active learning ($BMLP_{active}$) can reduce the requirement for training examples compared with passive learning (random experiment selection) to achieve a certain learning performance. In the context of scientific discovery, the experimental cost to arrive at a finding can be reduced by active learning since each actively selected experiment can have a higher return of information.
8.3 Experiment 3: learning digenic functions

### v4

@@ Appendix E Active learning sample complexity
According to the Blumer bound [7], for a version space learner that randomly samples examples to produce a hypothesis of error at most $\epsilon$ with a probability less than or equal to $\delta$, the sample complexity is $s_{passive}>\frac{1}{\epsilon}(ln(|H|)+ln(\frac{1}{\delta}))$. Following this, it must be true that $\epsilon\,s_{passive}>(\epsilon+\phi)\,s_{active}$. So we can derive the Inequality 1 in the theorem:
@@
Appendix F Implementation details
F.1 Encoding a metabolic network
Our encoding of a metabolic network is based on the Robot Scientist [34, 11]. The Robot Scientist is implemented using the ILP system Progol [45] and finds the producible set of metabolites by traversing a metabolic network. It uses a metabolic network as a deductive database and answers a query of user-defined gene deletion and experimental conditions based on SLD resolutions. Its deductive database is written in Prolog syntax and uses Prolog list operations to manipulate constant symbols representing metabolic substances. Metabolic reactions are enumerated based on the set of metabolites currently producible. The products of a metabolic reaction extend the list of producible metabolites. Since the number of metabolic reactions is finite, the closure of members in this list needs to contain essential metabolites for the growth of an organism. In an auxotrophic mutant experiment outcome prediction, a gene function is removed. Some key metabolic reaction pathways might no longer be viable due to the removal, leading to insufficient metabolites for growth. In this case, the predicted experiment’s outcome is a lack of growth and a phenotypic effect.
@@ F.1 Encoding a metabolic network
In the following program $\mathcal{P}_{3}$, the $codes/2$ clause says that the gene locus “b2379” is responsible for the enzyme “enz201”. The $enzyme/6$ clause defines the reversible reaction “alata_l” (L-alanine transaminase). This reaction involves reactants “akg_c” (Oxoglutarate), “ala__L_c” (L-alanine) and products “glu__L_c” (L-glutamate), “pyr_c” (Pyruvate). An integer value in {1,2} indicates the reaction reversibility and the integer “2” says that the reaction is reversible.
$\mathcal{P}_{3}:\left\{\begin{aligned} codes(b2379,enz201).&\\ enzyme(enz201,[b2379],[alata\_l],2,[[akg\_c,ala\_\_L\_c]],&\\ [[glu\_\_L\_c,pyr\_c]]).&\end{aligned}\right\}$
@@
We included this encoding in $BMLP_{active}$ for hypothesis generation and human interpretation. The $codes/2$ and $enzyme/6$ clauses enable us to describe the gene-reaction-metabolite mappings and standard identifiers in iML1515444The iML1515 model is described in the text-based JavaScript Object Notation (JSON). It uses standard identifiers from the Biochemical Genetic and Genomic (BiGG) [35] knowledge base of large-scale metabolic reconstructions. into a logic program containing $codes/2$ and $enzyme/6$ clauses. iML1515 uses standard nomenclature so the logic program encoding can be applied to other genome-scale metabolic networks. The logic program encoding can also be linked to publications and external databases to obtain further details on genes, proteins, reactions and metabolites.
F.2 BMLP-IE predictions from GEM
@@
Algorithm 3 shows the application of BMLP-IE (Algorithm 2) for finding the transitive closure of producible metabolites. The two Boolean matrices inputs to the BMLP-IE algorithm represent reactants and products participating in the reactions. The growth medium condition is encoded as the input vector to the BMLP-IE algorithm. This vector would be extended to represent metabolites that are producible by the organism in an auxotrophic mutant experiment. Specifying knockout genes excludes a subset of reactions. A vector computed during an intermediate step in the BMLP-IE algorithm denotes viable reactions given the set of knockout genes. A viable reaction is indicated by a binary bit 1 in this vector. Blocked reactions are excluded based on $codes/2$ and $enzyme/6$ definitions. In BMLP-IE, Boolean operations compute the transition between metabolic states and return the closure set of producible metabolites. A binary phenotypic classification is made according to the inclusion of essential metabolites in the closure set, indicating cell growth comparable to an unedited wild-type MG1655 E. coli strain. A binary classification of “1” corresponds to a phenotypic effect and “0” is no phenotypic effect. The essential metabolites come from the biomass function of the wild-type in iML1515 [43] [^5] We use the wild-type biomass function specified by iML1515 from http://bigg.ucsd.edu/ [35]. We use the qualitative version of this biomass function, meaning that we only care about the metabolites in the biomass, not their reaction coefficients.. The biomass function describes the necessary metabolite composition for the growth of the organism.
Appendix G Experimental details
@@
Since a logic program representing a GEM describes many biological entities and intricate relations between them, abduction of gene function hypotheses and in-vivo validations requires exploring a large hypothesis space and instance space. Both are challenging computationally and empirically. $BMLP_{active}$ addresses these by evaluating a large logic program with Boolean matrices and actively selecting experiment instances to perform. We demonstrated that the $BMLP_{active}$ framework can efficiently perform phenotypic predictions and actively suggest informative experiment instances that can reduce experiment costs.
G.1 Experimental cost
@@
Here, we defined the experimental cost function as the cost of the reagents, which are optional nutrient substances used in selected experiments. We also assumed that when added to growth media, reagents are used in the same volume. In addition, we did not explicitly account for plastic consumables, equipment deprecation or specify the time to carry out each experiment, which we consider fixed and are affected mainly by the number of experiments. We also did not account for costs incurred due to different experiment types, such as different measurement timescales that prevent experiments from being performed concurrently. To maintain consistent scaling for comparisons, the total experimental reagent cost was the amount of unit cost required by selected experiments. We considered the unit cost of the reagents as the cost of the cheapest reagent. We provided the costs of reagents in Table 2 and 3. We calculated the total experimental reagent cost as the sum of the reagent costs in selected experiments divided by the unit cost.
G.2 Experiment 1: BMLP-IE predictions
@@
Materials. We used $BMLP_{active}$ to predict a phenotype in each auxotrophic single-knockout mutant experiment. Every experiment removes one gene from the biological system and uses a growth medium. The GEM iML1515 has been validated against 16 carbon sources as optional nutrients in [43]. We focused on the same experimental conditions for base SWI-Prolog and $BMLP_{active}$. Experimental phenotypic data on the same 16 carbon sources and FBA phenotypic predictions of iML1515 are already processed as binary classes in [43]. In [43], measured and predicted mutant growth rates at a proportion of the wild-type’s growth rate greater than a fixed threshold are mapped to the no phenotypic effect class.
@@
Methods. We compared the runtime of $BMLP_{active}$ and the Robot Scientist [34, 11]’s encoding of the metabolic network (Appendix F). We first compared the mean and standard deviation in runtime of our new logical matrix encoding against the SWI-Prolog encoding. In addition, we explored multi-threading with both encodings. We used the multi-threading library offered by SWI-Prolog. Owing to the overhead of managing threads, multi-threading Boolean operations with SWI-Prolog have diminishing returns. In BMLP-IE, conditions are encoded as matrices, enabling us to predict in batches. For BMLP-IE, we used the concurrent threading library in SWI-Prolog at the batch level. The runtime experiments are repeated 10 times.
G.3 Experiment 2: active learning sample complexity and experiment cost
@@
Materials. Up to 3 of 16 common carbon sources were used as additional nutrients to the base growth medium. Each of these 16 carbon sources is an optional nutrient in the experimental data for validating [43]. This set contains preferred carbon sources [41, 76] for E. coli such as glucose and glycerol. We used iML1515 to predict with up to 3 optional carbon sources, generating $\sum_{0}^{i=3}\binom{16}{i}=697$ synthetic data. This allows us to go beyond the 16 experimental data points for each gene knockout in [43]. We observed a single-knockout phenotypic effect from 213 genes. We treated these synthetic data as ground truths, where 213 genes have only positive phenotype examples. We only focused on a subset of 3 genes because they have both positive and negative examples [^6] The distribution of training examples for this experiment is described in Table 4 of Appendix I.. Then, we removed gene functions associated with these genes from the background knowledge. Thus, the resulting incomplete model has a smaller set of $codes/2$ clauses.
@@
Methods. We constrained the selection of experiments by an experimental reagent cost budget. After experiment selection, both methods returned the hypothesis with the single highest compression, otherwise, a hypothesis was selected from those sharing the highest compression randomly. We focused on one removed gene function at a time to recover it. Each hypothesis, except the empty one, is associated with a gene locus identifier. We concentrate on 1295 candidate hypotheses that represent the non-maintenance reactions in iML1515 (including an empty hypothesis). The instance space had experiment candidates with optional growth nutrients up to 3 carbon sources, which was a total of 697 synthetic data points.
G.4 Experiment 3: learning digenic functions